THE ANALYSIS OF THE ROVIBRATIONAL SPECTRUM OF THE HO_{2} COMPLEX

\underline{XUDONG T. WU} and EDWARD F. HAYES, Dept. of Chemistry, The Ohio State University, Columbus, OH 43210.

An efficient algorithm has been developed that can be applied to full dimensional calculations of the rovibrational spectrum of triatom molecules. Performance tests of the algorithm on a massively parallel computer, such as the CRAY T3D, indicate that the algorithm scales almost linearly as the number of processors is increased from 4 to 256 processors. Using this parallel implementation of the algorithm on the CRAY T3D, we have calculated the rovibrational spectrum and wavefunctions of HO_{2} using the DMBE IV surface of Pastrana, Quintales, Brandao and Varandas.^{a} The performance of the algorithm for high values of the total angular momentum, J, is very good. For example, for J = 15 the algorithm achieves over 10 GFlops performance with 256 processors. With this level of performance, it is feasible to predict the spectrum of HO_{2} over a wide range of J values. The calculated spectrum will be compared with experimental results where available and with predictions made with simpler theoretical models.

a. M. R. Pastrana, L. A. M. Quintales, J. Brandão, and A. J. C. Varandas, J. Phys. Chem. 94, 8073 (1990).